An aircraft flying at an unknown initial velocity accelerates uniformly,at the rate of 5m/s².it reaches a velocity of 200m/s after accelerating over a distance of 3000 m. calculate the initial velocity of the aircraft

1
Expert's answer
2022-03-20T18:47:26-0400

Answer

Initial velocity

"u=\sqrt{v^2-2as}\\=\sqrt{200^2-2*5*300}\\=\sqrt{40000-3000}\\=192.35m/sec"